OpenStudy (anonymous):

identify the vertex, axis of symmetry and x-intercepts then check your results algebraically for g(x)=x^2+8x+11 (Show work please)

OpenStudy (anonymous):

anyone? Medal for a good answer:)

OpenStudy (radar):

Calculate the x intercepts using the quadratic formula which results in \[-4\pm \sqrt{5}\] That is a start.

OpenStudy (radar):

The function is already in standard form with a=1, and b=8 h=-b/2a h=-8/2=-4 substitute -4 for x in the function to obtain y (-4)^2 +8(-4)+11=16-32+11=-5 y=-5 vertex is at (-4,-5) Since the coefficient of x^2 is positive the parabola opens upward.

OpenStudy (radar):

I would assume the axis of symmetry would be the vertical line x=-4